OpenStudy (anonymous):

Given that logbase4(3)=a, and logbase4(5)=b express 375 in terms of a and/or b. there's nothing in 375? i tried 375logbase4(375)^(375) but it didn't work..

OpenStudy (anonymous):

@jim_thompson5910 you have any idea how to solve this? ^^

jimthompson5910 (jim_thompson5910):

just 375? no log with that?

OpenStudy (anonymous):

nope

jimthompson5910 (jim_thompson5910):

what is 4^a equal to?

OpenStudy (anonymous):

3?

jimthompson5910 (jim_thompson5910):

and 4^b ?

OpenStudy (anonymous):

5

jimthompson5910 (jim_thompson5910):

so that's sort of the stepping stone to getting the final answer

jimthompson5910 (jim_thompson5910):

notice how 375 = 15*25 = 3*5*5*5

OpenStudy (anonymous):

(4^b)^3 x 4^a?

jimthompson5910 (jim_thompson5910):

yes, so \[\Large 4^{3b}*4^a = 4^{3b+a} = 4^{a+3b}\]

OpenStudy (anonymous):

ahh thanks<3

jimthompson5910 (jim_thompson5910):

np